Kendra has $3.75 in dimes and nickels. If she has 2 times as many dimes as nickels, how many of each coin does she have?

Answer by TimothyLamb(4379)

5n + 10d = 375
d = 2n
---
5n + 10d = 375
2n - d = 0
---
solve the above system of linear equations for n and d:
---
answer:
n = 15
d = 30
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php